OpenStudy (anonymous):

Joe and his friend Thomas want to buy the same model skateboard. Joe needs $73 more and Thomas needs $65 more. If they combine their money, they will have just enough for ONE skateboard they can share. what is the cost of the skateboard? (It seems easy but I'm thrown)

Vocaloid (vocaloid):

hint: let x = cost of skateboard Joe = x - 73 Thomas = x - 65 Joe + Thomas = x solve for x

OpenStudy (anonymous):

I'm sorry I'm still confused? Wouldn't it be +?

Vocaloid (vocaloid):

why?

Vocaloid (vocaloid):

Joe and Thomas both have LESS than the skateboard, so you subtract

OpenStudy (anonymous):

oh okay well,the first hint you gave. do you mean x-75 = x-65?

Vocaloid (vocaloid):

nope.

Vocaloid (vocaloid):

Joe and Thomas have different amounts of money

Vocaloid (vocaloid):

Joe is x - 75 Thomas is x - 65 so you add them together and set them equal to x

Vocaloid (vocaloid):

*73

Vocaloid (vocaloid):

so... what is (x-65) + (x-73)?

OpenStudy (anonymous):

2x-138

OpenStudy (anonymous):

x= x-73+x-65 could this work? It comes out to 138

Vocaloid (vocaloid):

good!

OpenStudy (anonymous):

So thats it? lol thank you